# Scaling Law with Learning Rate Annealing

### v1

@@ 6.2 Possible Root Reasons of Delay Phenomenon in Learning Rate Annealing
No matter how LR changes, $S_{1}$ is always increasing over steps, resulting in consistently reducing the validation loss brought from $S_{1}$. Therefore, the forward area, $S_{1}$ would lengthen delay steps in LR annealing then re-warmup, but would shorten delay steps in LR re-warmup then annealing. The delay phenomenon is indeed related to $S_{2}$.
@@
The delay phenomenon could be intuitive in some cases. For example, suppose that learning rate decreases directly from 2e-4 to 2e-5 in one step, and maintains 2e-5. In this case, although the loss would decrease to a lower value but the parameter changes in one step is too small in neural networks. Given a sudden low LR, neural networks still require some steps to gradually optimize to a local minimum, incurring delay phenomenon. But still, analysis above still ends with a rough description, and we have not figured out the root reasons and look forward to completing this part in future work.
@@ 6.3 Other Possible Scaling Law Formats with LR annealing
$m_{i}$ $=\lambda\cdot m_{i-1}+(\eta_{k-1}-\eta_{k})$ (10)
$=\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k}$
$S_{2}$ $=\sum\limits_{i=1}^{s}m_{i}\color[rgb]{1,0,0}\cdot\eta_{i}^{\epsilon}$
@@
We have tried the revised function to fit data. We find that the fitting results are quite similar and $\epsilon$ is very close to 0, showing little use in practical effect. Hence, we adopt the original format in our experiments  [^8] We still recommend future developers to try this format if possible..
@@
However, the empirical results, as depicted in Fig. 18, demonstrate that the fitting improvement with the inclusion of $\zeta$ is quite marginal when compared to the version without this parameter. This slight enhancement does not justify the additional complexity introduced by managing negative values of $S_{2}$. Furthermore, the empirical observation that $\zeta$ tends to converge close to $1$ (e.g. $1.125$ in Fig. 18(c)) reinforces the notion that the original formulation of the function, without the power term $\zeta$, is adequately robust. This finding suggests that the direct influence of the learning rate annealing area, as initially modeled, sufficiently captures the essential dynamics without the need for this additional complexity  [^9] Another additional complexity lies in that $S_{2}^{\zeta}$ becomes incalculable when $S_{2}<0$ in LR re-warmup..
@@ 7.1 Tuning Scaling Law Format
We have tested a variety of equation formats to enhance the accuracy of the entire training process. As a result, the final equation format, as presented in Eq. 1, proves to be optimal so far across a range of scenarios. And the equation has achieved a level of practicality that enables the prediction of future loss when scaling training steps and model sizes. We will persist in refining the equation format to forecast the loss with the highest possible accuracy.
@@ 7.2 Extension to Post-training
In this work, we research primarily on the scope of pre-training of LLM. We also show how to apply our equation to guide the LR re-warmup strategy in continual pre-training. We will continue researching on how to extend our equation to post-training, which might include data distribution shift, data mixture, model alignment, and specific downstream evaluations.
@@ 8 Conclusion
In conclusion, we discover that the loss curves of neural language models empirically conform to a scaling law with learning rate annealing over training steps ($s$), which is expressed as $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$. This equation points out the art of balance between forward area and annealing area of learning rate schedulers, offering theoretical verification and explanation for numerous experimental findings of prior studies. We present the underlying theory of our equation including curve similarity and LR annealing momentum. We extend our equation to the format with model size $N$. Moreover, we prove that our equation is a generalized form of typical scaling laws, further showing the superiority of our equation.
Compared to typical scaling laws which predict the endpoint of one full loss curve, our equation can accurately forecast the loss of language model training at any given step and across any learning rate scheduler. As for required computation resources, our equation expends less than 1% of the computational cost required by typical scaling laws to fit language modeling loss. Our proposed approach greatly democratizes the fitting and prediction of scaling laws in the development of large language models.
@@ Appendix A Impact of Warmup Steps
We conduct experiments on the imapct of learning rate warmup steps. As shown in Fig. 19, we find that 500 warmup steps can get the lowest validation loss compared to 100 or no warmup steps. The experimental results also guide us to choose 500 warmup steps in our experiments of this work.
Figure 19: The comparison of the true loss curve of different warmup steps. We experiment on cosine LRS with 20K total steps.
@@ Appendix B Experimental Setups
In this work, we use multiple sets of experimental setups, in order to validate that our equation can work across different experimental setups. For clarification, we present the experimental setup list as shown in Table 3.
In our most experiments (including our takeaway sections), we use the main setting A. In our fitting and prediction experiment, we repeat experiments and consolidate our conclusion using another setting, setting B, and the experimental results are shown in Fig. 21. In our $N$-extended scaling law fitting experiment, we repeat experiments and consolidate our conclusion using another setting, setting C, and the experimental results are shown in Fig. 22.
Setups Setting A (mainly) Setting B Setting C
Model Size (Non-embedding) $594$M $293$M multiple
@@
Maximal Learning Rate $2\times 10^{-4}$ $2\times 10^{-4}$ $1.381\times 10^{-3}$
@@
Tokenizer LLAMA-3’s LLAMA-3’s Extended LLAMA-2’s
$\beta_{1}$,$\beta_{2}$ in Adam Optimizer $0.9$, 0.90.90.90.95 $0.9$, $0.95$ $0.9$, $0.95$
@@
Table 3: Experimental settings adopted in this work. * denotes pre-training multilingual dataset including mixture of sources such as common crawls, books, arxiv, code, etc. “Extended LLAMA-2’s” is extended from LLAMA-2’s tokenizer  (Meta, 2023) by adding vocabulary.
Appendix C Our Scaling Law on Another Experiments Setups
Fig. 2 and Fig. 3 show that our equation can work very well on our main experimental setup. For proving that our scaling law with LR annealing can apply to different (but given) experimental settings, we change the setting from $A$ to $B$ (refer to Table 3) and observe whether our equation can still work or not. The fitting results are shown in Fig. 20. The prediction results are shown in Fig. 21. The results suggest that our scaling law with LR annealing can still work well across different experimental setups.
@@ Appendix C Our Scaling Law on Another Experiments Setups
Figure 20: Full loss curve fitting on cosine (30K steps to $\eta_{min}=0$) and constant LRS. The figures omit the warmup in the first 100 steps. After fitting, we get a universal loss equation $L=2.761+0.517\cdot S_{1}^{-0.491}-0.458\cdot S_{2}$. Refer to setting B in Table 3 for experimental setups.
@@
Figure 21: Full loss curve prediction (120K steps) by the universal loss curve equation across various LRS. The left, the medium, and the right figures in each row are learning rate curve, zoomed-out loss prediction, and zoomed-in loss prediction, respectively. The red rectangle means the zoomed-in zone. The figures omit the warmup in the first 100 steps. The universal loss curve is $L=2.761+0.517\cdot S_{1}^{-0.491}-0.458\cdot S_{2}$, fitted as in Fig. 3(b). The equation can accurately predict various forms of unseen LRS, and it is highly precise in predicting the trend of loss changes as LR varies. Please note that these are predictive results, which means that none of the points in this figure (except constant LRS) are involved in the fitting process. The mean prediction errors across various LRS are low to $\sim 0.2\%$. Refer to setting B in Table 3 for experimental setups.
Appendix D Our $N$-extended Scaling Law on Another Experiments Setups
@@
Fig. 6(b) show that our $N$-extended equation can work very well on our main experimental setup. Similarly, for proving that our $N$-extendec caling law can apply to different (but given) experimental settings, we change the setting from $A$ to $C$ (refer to Table 3) and observe whether our equation can still work or not. The fitting results are shown in Fig. 22. The results suggest that our $N$-extended scaling law with LR annealing can still work well across different experimental setups.
@@
Figure 22: Curve fitting on cosine LRS (143K steps to $\eta_{min}=0$) of many model sizes using our scaling law extended to model size $N$. Refer to setting C in Table 3 for experimental setups.
@@
Appendix E 1-sqrt and 1-square Annealing
The 1-sqrt annealing is proposed by Hägele et al. (2024), whose learning rate at step $s$ in annealing stage is defined as:
@@
$\eta_{s}=\eta_{min}+f(s)\cdot(\eta_{max}-\eta_{min})$
We draw the learning rate curve of WSD (20% and 50% 1-sqrt annealing) in Fig. 23, compared with cosine annealing. Other than 1-sqrt annealing, Hägele et al. (2024) also mentions 1-square annealing method defined as:
@@
$\eta_{s}=\eta_{min}+f(s)\cdot(\eta_{max}-\eta_{min})$

### v2

@@ 6.2 Possible Root Reasons of Delay Phenomenon in Learning Rate Annealing
No matter how LR changes, $S_{1}$ is always increasing over steps, resulting in consistently reducing the validation loss brought from $S_{1}$. Therefore, the forward area, $S_{1}$ would lengthen delay steps in LR annealing then re-warmup, but would shorten delay steps in LR re-warmup then annealing. The delay phenomenon is indeed related to $S_{1}$.
@@
The delay phenomenon could be intuitive in some cases. For example, suppose that learning rate decreases directly from 2e-4 to 2e-5 in one step, and then maintains 2e-5. In this case, although the loss would decrease to a lower value but the parameter changes in one step is too small in neural networks. Given a sudden low LR, neural networks still require some steps to gradually optimize to a local minimum, incurring delay phenomenon. But still, analysis above still ends with a rough description, and we have not figured out the root reasons and look forward to completing this part in future work.
@@ 6.3 Other Possible Scaling Law Formats with LR annealing
$m_{i}$ $=\lambda\cdot m_{i-1}+(\eta_{i-1}-\eta_{i})$ (9)
$=\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k},$
$S_{2}$ $=\sum\limits_{i=1}^{s}m_{i}\color[rgb]{1,0,0}\cdot\eta_{i}^{\epsilon},$
@@
We have tried the revised function to fit data. We find that the fitting results are quite similar and $\epsilon$ is very close to 0, showing little use in practical effect. Hence, we adopt the original format in our experiments. However, we still recommend future developers to try this format if possible.
@@
However, the empirical results, as depicted in Fig. 19, demonstrate that the fitting improvement with the inclusion of $\zeta$ is quite marginal when compared to the version without this parameter. This slight enhancement does not justify the additional complexity introduced by managing negative values of $S_{2}$. Furthermore, the empirical observation that $\zeta$ tends to converge close to $1$ (e.g. $1.125$ in Fig. 19(c)) reinforces the notion that the original formulation of the function, without the power term $\zeta$, is adequately robust. This finding suggests that the direct influence of the learning rate annealing area, as initially modeled, sufficiently captures the essential dynamics without the need for this additional complexity. Another additional complexity lies in that $S_{2}^{\zeta}$ becomes incalculable when $S_{2}<0$ in LR re-warmup.
@@ 7.1 Tuning Scaling Law Format
We have tested a variety of equation formats to enhance the accuracy of the entire training process. As a result, the final equation format, as presented in Eq. 1, proves to be optimal so far across a range of scenarios. We add only one extra parameter but obtain a very good fitting and prediction result. The formulation has achieved a level of practicality that enables the prediction of future loss when scaling training steps and model sizes. We expect more following researches to explore the format of the scaling law with learning rate annealing.
7.2 More Applications via Our Scaling Law
In Sec. 4, we present many instances to apply our scaling law with LR annealing, to predict future training dynamics with a cost-free manner. We believe that our equation can help analyze and select more training recipes in specific scenarios.
@@ 7.3 Extension to Post-training
In this work, we research primarily on the scope of pretraining of LLM. We also show how to apply our equation to guide the LR re-warmup strategy in continual pre-training. We will continue researching on how to extend our equation to post-training, which might include data distribution shift, data mixture, model alignment, and specific downstream evaluations.
@@ 8 Conclusion
In conclusion, we propose that the loss curves of neural language models empirically adhere to a scaling law with learning rate annealing over training steps $s$: $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$. This equation can accurately predict full loss curves across unseen learning rate schedulers. We present the underlying intuition and theory for deriving our equation and demonstrate that our approach can be extended to capture the scaling effect of model sizes. Extensive experiments demonstrate that our proposed scaling law has good accuracy, scalability, and holds under various experimental setups. It also offers accurate theoretical insights to the training dynamics of LLMs, and explains numerous phenomena observed in previous studies. We believe that the scaling law with LR annealing is promising to reshape the understanding of researchers for LLM training and scaling laws.
@@ Appendix A Impact of Warmup Steps
We investigate the impact of LR warmup steps on loss curves. Specifically, different values of warmup steps are experimented. As shown in Fig. 20, we find that a warmup step value of 500 accelerates convergence and achieves the lowest validation loss compared to 100 or no warmup. This finding is aligned with previous works Liu et al. (2020); Kosson et al. (2024).
Based on these findings, we use 500 warmup steps in our main experiments when training models from scratch. Note that LR warmup when training a randomly initialized model is different from LR re-warmup in continual training. In fact, the re-warmup process can be seen as a special types of LRS where $S_{2}<0$ in our equation.
Figure 20: Validation loss curves corresponding to different warmup steps. We use cosine LRS with 20K total steps.
@@ Appendix B Experimental Setups
In this work, we use multiple experimental setups to validate the effectiveness of our equation across a variety of conditions. For clarity, we summarize all experimental setups in Table 3. The majority of our experiments use Setting A. Additionally we also successfully fit our equation to the loss curves of BLOOM (BigScience, 2022) and OLMo (Groeneveld et al., 2024). The experimental settings used to produce these loss curves are significantly different from ours. Please refer to their papers for details on their specific experimental setups.
Table 3: Experimental settings adopted in this work. Model size denotes the number of non-embedding parameters. Our datasets include Fineweb (Penedo et al., 2024) and RedPajama-CC (Computer, 2023). * denotes pre-training multilingual datasets including mixture of sources such as common crawls, books, arxiv, code, etc. We use AdamW Optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2017), denoted as AO. Most experiments adopt Llama-3’s tokenizer (Dubey et al., 2024). Ext Llama-2’s is extended from Llama-2’s tokenizer  (Touvron et al., 2023) by adding vocabulary.
Setups Setting A (main) Setting B Setting C
Model Size $594$M $293$M multiple
@@
Maximal LR $2\times 10^{-4}$ $2\times 10^{-4}$ $1.381\times 10^{-3}$
@@
Sequence Length $4096$ $4096$ $4096$
Tokenizer Llama-3’s Llama-3’s Ext Llama-2’s
$\beta_{1}$,$\beta_{2}$ in AO $0.9$, 0.90.90.90.95 $0.9$, $0.95$ $0.9$, $0.95$
@@
Setups Setting D (MoE) Setting E (1.4T tokens)
Model Size $8\times 106$M $1704$M
Train Dataset Fineweb Mixture-train*
Val Dataset RedPajama-CC Mixture-valid*
Total Steps $60$K $350$K
Maximal LR $2\times 10^{-4}$ $6\times 10^{-4}$
Warmup Steps $500$ $1000$
Batch Size (tokens) $4$M $4$M
Sequence Length $4096$ $8192$
Tokenizer Llama-3’s Llama-3’s
$\beta_{1}$,$\beta_{2}$ in AO $0.9$, 0.90.90.90.95 $0.9$, $0.95$
Top-$k$ Experts $2$
Auxiliary Loss $0.01$
Weight Decay $0.1$ $0.1$
Gradient Clip $1.0$ $1.0$
Appendix C Results on Extensive Experiments Setups
C.1 Another Set of Training Hyper-parameters
Fig. 2 and Fig. 3 show that our equation can work very well on our main experimental setup detailed in Table 3. To validate the effectiveness of our scaling law formulation on different experimental settings, we report the results on Setting $B$ (see Table 3). The fitting results are shown in Fig. 21, and the prediction results are shown in Fig. 22. The results suggest that our scaling law with LR annealing works well across different experimental setups.
@@ C.1 Another Set of Training Hyper-parameters
Figure 21: Using Eq.1 to fit full loss curves yield by cosine and constant LRS. Total steps=30K, $\eta_{min}=2\times 10^{-4}$, $\eta_{min}=0$. We omit the warmup steps in the figure. The fitted equation is $L=2.761+0.517\cdot S_{1}^{-0.491}-0.458\cdot S_{2}$. Refer to setting B in Table 3 for detailed experimental setups.
@@
Figure 22: Using the fitted equation in Fig. 21 to predict full loss curves for unseen LRS with 120K total steps. The left, medium, and right columns represent LR curve, the predicted loss curves, and a zoomed-in view of loss curve, respectively. The red rectangle indicates the zoomed-in zone. Warmup steps (100) are not shown in this figures. Notable, all LRS and loss curves shown here were unseen during the fitting process. The mean prediction errors across different LRS is as low as $\sim 0.2\%$. Refer to setting B in Table 3 for detailed experimental setups.
C.2 Experiments on the Mixture of Experts Architecture
We validate our equation on the Mixture of Experts (MoE) architecture. The experimental setting is shown as setting $D$ in Table 3. We add the widely-used auxiliary loss for load balancing among experts (Fedus et al., 2021). Moreover, we change the LRS and total steps to 60K WSD with 10K annealing steps in fitting, testing whether our scaling law is effective under various circumstances. The fitting results shown in Fig. 23 and the prediction results shown in Fig. 24 indicates that our equation still works well on the MoE architecture.
(a) Learning rate.
(b) Zoomed-out loss fitting.
(c) Zoomed-in loss fitting.
Figure 23: Full loss curve fitting for MoE models. After fitting, we get a universal loss equation $L=2.801+0.424\cdot S_{1}^{-0.619}-0.347\cdot S_{2}$. Refer to setting D in Table 3 for detailed experimental setups.
(a) Full curve prediction of multi-step cosine LRS (80% + 10% + 10%) (DeepSeek-AI, 2024)
(b) Full curve prediction of WSD LRS (17% exponential annealing to $\eta_{min}=0$(Hu et al., 2024).
(c) Full curve prediction of Cyclic LRS (Smith, 2017) including multiple unseen LR re-warmup stages.
Figure 24: Using the fitted equation in Fig. 23 to predict full loss curves for unseen LRS (MoE model). The left, medium, and right columns represent LR curve, the predicted loss curves and a zoomed-in view of loss curves, respectively. The red rectangle indicates the zoomed-in zone. Warmup steps (500) are not shown in this figure. Noteable, all LRS and loss curves shown here were unseen during the fitting process. The mean prediction errors across different LRS is as low as $\sim 0.2\%$. Refer to setting D in Table 3 for detailed experimental setups.
C.3 Scaling Up for Much Longer Steps
In this section, we validate the effectiveness of our equation in predicting loss curves over significantly long total steps. This scalability is particularly useful in large-scale training. Specifically, we fit our equation on loss curves yield by the constant and cosine LRS with 30K steps and predict the loss curve for the WSD LRS with 350K steps. The model has 1.7 billion parameters, and the training uses 1,400 billion tokens. More detailed experimental setup is shown as the setting $E$ in Table 3.
The fitting and prediction results are shown in Fig. 25 and Fig. 26 respectively. It shows that we accurately predict the loss curve for the annealing stage after 10x longer steps in advance.
(a) Learning rate.
(b) Zoomed-out loss fitting.
(c) Zoomed-in loss fitting.
Figure 25: Full loss curve fitting for the constant and cosin LRS on 30K steps. After fitting, we get a universal loss equation $L=2.788+0.906\cdot S_{1}^{-0.416}-0.254\cdot S_{2}$. Refer to setting E in Table 3 for detailed experimental setups.
(a) Learning rate.
(b) Zoomed-out loss prediction.
(c) Zoomed-in loss prediction.
Figure 26: Using the fitted equation in Fig. 25 to predict the full loss curve equation under the WSD LRS (10% cosine annealing ratio to $\eta_{min}=0$). Our equation accurately predict the loss curve in the annealing stage after the 10x longer steps. Refer to setting E in Table 3 for detailed experimental setups.
C.4 Open-sourced Models
To further validate our proposed scaling law, we apply our equation on modeling the loss curves of open-sourced language models, including BLOOM-176B (BigScience, 2022) and OLMo-1B (Groeneveld et al., 2024). As shown in Fig. 27, our equation fits very well on loss curves of open-sourced models, even when the model size scales up to 176B (e.g. BLOOM) and token number scales up to 2000B over 740K steps (e.g. OLMo).
(a) Full loss curve fitting on BLOOM-176B.
(b) Full loss curve fitting on OLMo-1B (2T tokens).
Figure 27: Fitting loss curves for open sourced models. We extract the curve of BLOOM from https://huggingface.co/bigscience/bloom/tensorboard, and choose the column lm-loss-validation/valid/lm loss validation as the validation loss. We extract the curve of OLMo from https://wandb.ai/ai2-llm/OLMo-1B?nw=nwuserdirkgr, and choose the column eval/pile/CrossEntropyLoss as the validation loss. Both models adopt cosine LRS.
C.5 Our $N$-extended Scaling Law on Another Experiments Setups
@@
Fig. 7(b) shows that our $N$-extended equation (Eq. 7) works well on our main experimental setup. Similarly, to validate the effectiveness of our $N$-extended scaling equation on various different experimental settings, we change our setup from setting $A$ to setting $C$ (refer to Table 3). The fitting results are shown in Fig. 28. The results suggest that our $N$-extended scaling law with LR annealing works well across different experimental setups.
@@
Figure 28: Curve fitting on cosine LRS (143K steps to $\eta_{min}=0$) for various different model sizes using our scaling law extended to model size $N$. Refer to setting C in Table 3 for detailed experimental setups.
Appendix D WSD Scheduler and Annealing Functions
@@
$WSD(s)=\begin{cases}&\frac{s}{T_{\text{warmup}}}\eta_{max},\quad s\leq T_{\text{warmup}}\\ &\eta_{max},\quad T_{\text{warmup}}<s\leq T_{\text{stable}}\\ &\eta_{min}+f(s)\cdot(\eta_{max}-\eta_{min}),\quad T_{\text{stable}}<s\leq T_{\text{total}}\end{cases}$ (10)
Where $0\leq f(s)\leq 1$ is typically a decreasing function about step $s$, and $\eta_{max}$ is the maximal learning rate. Hägele et al. (2024) consolidate the effectiveness of the WSD scheduler with numerious empirical experiments. Moreover, Hägele et al. (2024) also find that using 1-sqrt annealing and a moderate annealing ratio (e.g. 20%) can further decrease the final loss. The 1-sqrt annealing is defined as:
@@
Also,  Hägele et al. (2024) mention the 1-square annealing method as a baseline, which is defined as:
@@
We draw the LR curve of the WSD LRS proposed by Hu et al. (2024) (20% and 50% 1-sqrt annealing) along with the WSD LRS with a cosine annealing in Fig. 29.